a car traveling with initial velocity of 22.4 m/s skids to a stop in 22.5 displacement. What is the skidding displacement of the car


1
Expert's answer
2021-08-13T10:20:41-0400

Given:

"v_i=22.4\:{\rm m/s},\quad v_f=0."

"d=22.5\:\rm m"

The displacement

"d=\frac{v_f^2-v_i^2}{2a}"

Hence, the acceleration

"a=\frac{v_f^2-v_i^2}{2d}=\frac{0^2-22.4^2}{2*22.5}=-11.2\:\rm m/s^2"